# Dual-Source SPIR over a noiseless MAC without Data Replication or Shared Randomness

### v1

@@ IV-A Special case $L_{1}=L_{2}=2$
Suppose that $L_{1}=L_{2}=2$. Our coding scheme is described in Algorithm 1 and analyzed as follows.
@@
2:The servers use the channel (1) as follows:
@@
4:b: Servers $1$ and $2$ send $X_{1}^{n}$ and $X_{2}^{n}$, respectively, over the channel (1)
@@
$\mathcal{G}\triangleq\{i\in\llbracket 1,n\rrbracket:Y_{i}\in\{0,2\}\},\quad\mathcal{B}\triangleq\{i\in\llbracket 1,n\rrbracket:Y_{i}=1\};$
@@
10:d: Constructs $\mathcal{B}_{1},\mathcal{B}_{2}$ such that $\mathcal{B}_{1}\cup\mathcal{B}_{2}\subset\mathcal{B}$, $\mathcal{B}_{1}\cap\mathcal{B}_{2}=\emptyset$, and $(|\mathcal{B}_{1}|,|\mathcal{B}_{2}|)=(\alpha M,\bar{\alpha}M)$;
@@
Note that when $Z_{j}=i$, $i\in\{1,2\}$, $j\in\{1,2\}$, the client can determine $X_{j}^{n}[S_{i}^{(j)}]$.
@@
14:a: Server $1$ sends to the client $(M_{11},M_{12})$ where
@@
15:b: Server $2$ sends to the client $(M_{21},M_{22})$ where
@@ IV-A2 Privacy of the receiver selection
=$\stackrel{{\scriptstyle(a)}}{{=}}I(F_{1,1}F_{1,2}X_{1}^{n}M_{21}M_{22}S_{1}^{(1)}S_{2}^{(1)}S_{1}^{(2)}S_{2}^{(2)};Z_{1}Z_{2})$
=$\stackrel{{\scriptstyle(b)}}{{=}}I(F_{1,1}F_{1,2}X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)}S_{1}^{(2)}S_{2}^{(2)};Z_{1}Z_{2})$
=$\stackrel{{\scriptstyle(c)}}{{=}}I(X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)}S_{1}^{(2)}S_{2}^{(2)};Z_{1}Z_{2})$
=$\stackrel{{\scriptstyle(d)}}{{=}}I(X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)};Z_{1})+I(X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)};Z_{2}|Z_{1})$
$\phantom{--}+I(S_{1}^{(2)}S_{2}^{(2)};Z_{2}|X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)})+I(S_{1}^{(2)}S_{2}^{(2)};Z_{1}|Z_{2}X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)})$
$\stackrel{{\scriptstyle(e)}}{{\leqslant}}I(X_{1}^{n}[\mathcal{G}_{1}\cup\mathcal{B}_{1}]S_{1}^{(1)}S_{2}^{(1)};Z_{1})+I(X_{1}^{n}[\mathcal{G}_{2}\cup\mathcal{B}_{2}]S_{1}^{(2)}S_{2}^{(2)};Z_{2})$
@@
    holds because $I(X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)};Z_{2}|Z_{1})=0$ by independence between $Z_{2}$ and $(X_{1}^{n},S_{1}^{(1)},S_{2}^{(1)},Z_{1})$, $I(S_{1}^{(2)}S_{2}^{(2)};Z_{1}|Z_{2}X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)})\leqslant I(S_{1}^{(2)}S_{2}^{(2)}Z_{2}X_{1}^{n}[\mathcal{G}_{2}\cup\mathcal{B}_{2}];Z_{1}X_{1}^{n}[\mathcal{G}_{1}\cup\mathcal{B}_{1}]S_{1}^{(1)}S_{2}^{(1)})=0$, and we also have $I(S_{1}^{(2)}S_{2}^{(2)};Z_{2}|X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)})\leqslant I(S_{1}^{(2)}S_{2}^{(2)}X_{1}^{n}[\mathcal{G}_{2}\cup\mathcal{B}_{2}];Z_{2}|X_{1}^{n}[\mathcal{G}_{1}\cup\mathcal{B}_{1}]S_{1}^{(1)}S_{2}^{(1)})=I(S_{1}^{(2)}S_{2}^{(2)}X_{1}^{n}[\mathcal{G}_{2}\cup\mathcal{B}_{2}];Z_{2})$.
@@ IV-A3 Privacy of the unselected strings for the senders
We now prove that the client does not learn any information about the unselected files as follows
$I\left(Z_{1}Z_{2}Y^{n}\mathbf{A};F_{1,\bar{Z}_{1}}F_{2,\bar{Z}_{2}}\right)$
=$\stackrel{{\scriptstyle(a)}}{{=}}I\left(Z_{1}Z_{2}Y^{n}M_{11}M_{12}M_{21}M_{22};F_{1,\bar{Z}_{1}}F_{2,\bar{Z}_{2}}\right)$
=$\stackrel{{\scriptstyle(b)}}{{=}}\frac{1}{4}\sum_{z_{1},z_{2}}I\left(Y^{n}M_{11}M_{12}M_{21}M_{22};F_{1,\bar{z}_{1}}F_{2,\bar{z}_{2}}|Z_{1}=z_{1},Z_{2}=z_{2}\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}\frac{1}{4}\sum_{z_{1},z_{2}}I\left(Y^{n}F_{1,z_{1}}F_{2,z_{2}},X_{1}^{n}[S^{(1)}_{\bar{z}_{1}}]\oplus F_{1,\bar{z}_{1}},X_{2}^{n}[S^{(2)}_{\bar{z}_{2}}]\oplus F_{2,\bar{z}_{2}};F_{1,\bar{z}_{1}}F_{2,\bar{z}_{2}}|Z_{1}=z_{1},Z_{2}=z_{2}\right)$
=$\stackrel{{\scriptstyle(d)}}{{=}}\frac{1}{4}\sum_{z_{1},z_{2}}I\left(Y^{n},X_{1}^{n}[S^{(1)}_{\bar{z}_{1}}]\oplus F_{1,\bar{z}_{1}},X_{2}^{n}[S^{(2)}_{\bar{z}_{2}}]\oplus F_{2,\bar{z}_{2}};F_{1,\bar{z}_{1}}F_{2,\bar{z}_{2}}|Z_{1}=z_{1},Z_{2}=z_{2}\right)$
=$\stackrel{{\scriptstyle(e)}}{{=}}0,$
@@
    holds by (28) in Lemma 9 (Appendix A) with $\epsilon=\delta=0$ and the substitutions $A\leftarrow Y^{n}$, $C\leftarrow(F_{1,\bar{z}_{1}},F_{2,\bar{z}_{2}})$, $D\leftarrow(X_{1}^{n}[S^{(1)}_{\bar{z}_{1}}],X_{2}^{n}[S^{(2)}_{\bar{z}_{2}}])$.
@@ IV-A4 Servers’ mutual privacy
$I\left(F_{1,1}F_{1,2}X_{1}^{n}\mathbf{A};F_{2,1}F_{2,2}\right)$ =$\stackrel{{\scriptstyle(a)}}{{=}}I(F_{1,1}F_{1,2}X_{1}^{n}M_{21}M_{22}S_{1}^{(1)}S_{2}^{(1)}S_{1}^{(2)}S_{2}^{(2)};F_{2,1}F_{2,2})$
=$\stackrel{{\scriptstyle(b)}}{{=}}0,$
@@ IV-B Case $L_{1}=L_{2}$
In this section, we present an achievability scheme for the case $L_{1}=L_{2}=L$ separately, as it is significantly simpler than the achievability scheme for the general case $L_{1}\neq L_{2}$, which is discussed in Section IV-C. The idea is to construct a coding scheme for this case by utilizing $L-1$ times Algorithm 2 developed for the case $L_{1}=L_{2}=2$. This reduction idea is inspired by [18] in the context of oblivious transfer. Before we describe our coding scheme in Algorithm 2, we illustrate the coding scheme idea in the following example.
@@
Suppose that $L_{1}=L_{2}=3$, and $S$ and $T$ are uniformly distributed over $\llbracket 1,2^{nR_{1}}\rrbracket$ and $\llbracket 1,2^{nR_{2}}\rrbracket$, respectively. Servers 1 and 2 engage in a dual-source SPIR protocol from Section IV, interacting with the client twice. In the first interaction, the client selects between $S$ and $F_{1,1}$ from Server 1, and between $T$ and $F_{2,1}$ from Server 2. During the second interaction, the client chooses between $F_{1,2}\oplus S$ and $F_{1,3}\oplus S$ from Server 1, and between $F_{2,2}\oplus T$ and $F_{2,3}\oplus T$ from Server 2. Specifically, in the first round, if the client desires $F_{1,1}$, they request $F_{1,1}$ from Server 1, otherwise, they request $S$. In the second round, the client requests $F_{1,2}\oplus S$ or $F_{1,3}\oplus S$ from Server 1, depending on whether they wish to retrieve $F_{1,2}$ or $F_{1,3}$, respectively. The same logic applies to Server 2.
Algorithm 2 Dual-source SPIR when $L_{1}=L_{2}$
1:$L-2$ sequences $(S_{1,t})_{t\in\llbracket 1,L-2\rrbracket}$ uniformly distributed over $\{0,1\}^{nR_{1}}$, $L-2$ sequences $(S_{2,t})_{t\in\llbracket 1,L-2\rrbracket}$ uniformly distributed over $\{0,1\}^{nR_{2}}$, the file selection $(Z_{1},Z_{2})\in\mathcal{L}_{1}\times\mathcal{L}_{2}$
@@
$(C_{j,t}[1],C_{j,t}[2])$ $\triangleq(F_{j,t}\oplus S_{j,t-1},S_{j,t-1}\oplus S_{j,t}),\forall t\in\llbracket 2,L-2\rrbracket$
@@
For $t\in\llbracket 1,L-1\rrbracket$, define $C_{j,t}\triangleq(C_{j,t}[1],C_{j,t}[2])$.
@@
5:     The client and the two servers perform the SPIR protocol in Algorithm 1 with the two sequences $(C_{j,t}[1],C_{j,t}[2])$ at Server $j\in\llbracket 1,2\rrbracket$ and the selection $(Z_{1,t},Z_{2,t})$ for the client. The subscript $t$ is used in the notation of the random variables $(X_{1,t}^{n},X_{2,t}^{n},Y_{t}^{n},\mathbf{A}_{t},Z_{1,t},Z_{2,t})$ involved in this SPIR protocol.
@@
From Section IV-A5, the lengths of the files in Algorithm 1 for the case $L_{1}=L_{2}=2$ and $n$ channel uses is $\left(\alpha M,\bar{\alpha}M\right)$, hence, the rate pair achieved with Algorithm 2, which utilizes $L-1$ times Algorithm 1 is

### v2

@@ IV-A Special case $L_{1}=L_{2}=2$
Suppose that $L_{1}=L_{2}=2$. Our coding scheme, based on [ahlswede2013oblivious], is described in Algorithm 1. The idea of Algorithm 1 is to exploit the additive channel structure $Y^{n}=X_{1}^{n}+X_{2}^{n}$. Positions where $Y_{t}\in\{0,2\}$ (“good” indices) let the client recover the channel inputs, while positions with $Y_{t}=1$ (“bad” indices) reveal nothing about the inputs. Using this partition, the client constructs two sets of indices $(S_{1}^{(i)},S_{2}^{(i)})$ for each Server $i\in\{1,2\}$, one corresponding to recoverable positions, the other to unrecoverable ones, without revealing which is which. Each server then applies a one-time pad to its inputs over both sets as follows. On the set the client can recover, the masked input allows reconstruction of the desired file. On the set the client cannot recover, the masked input hides information about the non-requested file. This ensures the client obtains exactly the requested file while preserving both client and server privacy. Algorithm 1 is formally analyzed as follows.
@@
2:The channel (1) is used as follows:
@@
4:b: Servers $1$ and 22 send $X_{1}^{n}$ and $X_{2}^{n}$, respectively, over the channel (1)
@@
$\mathcal{G}$ $\triangleq\{i\in\llbracket 1,n\rrbracket:Y_{i}\in\{0,2\}\},$
$\mathcal{B}$ $\triangleq\{i\in\llbracket 1,n\rrbracket:Y_{i}=1\};$
@@
10:d: Constructs $\mathcal{B}_{1},\mathcal{B}_{2}$ such that $\mathcal{B}_{1}\cup\mathcal{B}_{2}\subset\mathcal{B}$, $\mathcal{B}_{1}\cap\mathcal{B}_{2}=\emptyset$, and $(|\mathcal{B}_{1}|,|\mathcal{B}_{2}|)=(\alpha M,\bar{\alpha}M)$; [^1] As an example, suppose $n=12$, $\alpha=\scriptstyle\frac{1}{3}$, and $Y^{n}=(1,0,2,0,1,2,0,1,1,2,1,1)$. Then, $M=6$, $\mathcal{G}=\{2,3,4,6,7,10\}$, $\mathcal{B}=\{1,5,8,9,11,12\}$, and, for instance, $\mathcal{G}_{1}=\{2,3\}$, $\mathcal{G}_{2}=\{4,6,7,10\}$, $\mathcal{B}_{1}=\{1,5\}$, $\mathcal{B}_{2}=\{8,9,11,12\}$.
@@
Note that when $Z_{i}=j$, $i\in\{1,2\}$, $j\in\{1,2\}$, the client can determine $X_{i}^{n}[S_{j}^{(i)}]$.
@@
14:a: Server 11 sends to the client $(M_{11},M_{12})$ where
@@
15:b: Server 22 sends to the client $(M_{21},M_{22})$ where
@@ IV-A2 Privacy of the receiver selection
$\stackrel{{\scriptstyle(a)}}{{=}}I(F_{1,1}F_{1,2}X_{1}^{n}M_{21}M_{22}S_{1}^{(1)}S_{2}^{(1)}S_{1}^{(2)}S_{2}^{(2)};Z_{1}Z_{2})$
$\stackrel{{\scriptstyle(b)}}{{\leqslant}}I(F_{1,1}F_{1,2}X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)}S_{1}^{(2)}S_{2}^{(2)};Z_{1}Z_{2})$
$\stackrel{{\scriptstyle(c)}}{{=}}I(X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)}S_{1}^{(2)}S_{2}^{(2)};Z_{1}Z_{2})$
$\stackrel{{\scriptstyle(d)}}{{=}}I(X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)};Z_{1})+I(X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)};Z_{2}|Z_{1})$
$\phantom{--}+I(S_{1}^{(2)}S_{2}^{(2)};Z_{2}|X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)})$
$\phantom{--}+I(S_{1}^{(2)}S_{2}^{(2)};Z_{1}|Z_{2}X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)})$
$\stackrel{{\scriptstyle(e)}}{{\leqslant}}I(X_{1}^{n}[\mathcal{G}_{1}\cup\mathcal{B}_{1}]S_{1}^{(1)}S_{2}^{(1)};Z_{1})$
$\phantom{--}+I(X_{1}^{n}[\mathcal{G}_{2}\cup\mathcal{B}_{2}]S_{1}^{(2)}S_{2}^{(2)};Z_{2})$
@@
    holds because $I(X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)};Z_{2}|Z_{1})=0$ by independence between $Z_{2}$ and $(X_{1}^{n},S_{1}^{(1)},S_{2}^{(1)},Z_{1})$,
    $I(S_{1}^{(2)}S_{2}^{(2)};Z_{1}|Z_{2}X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)})$
    $\leqslant I(S_{1}^{(2)}S_{2}^{(2)}Z_{2}X_{1}^{n}[\mathcal{G}_{2}\cup\mathcal{B}_{2}];Z_{1}X_{1}^{n}[\mathcal{G}_{1}\cup\mathcal{B}_{1}]S_{1}^{(1)}S_{2}^{(1)})$
    $=0,$
    and we also have
    $I(S_{1}^{(2)}S_{2}^{(2)};Z_{2}|X_{1}^{n}S_{1}^{(1)}S_{2}^{(1)})$
    $\leqslant I(S_{1}^{(2)}S_{2}^{(2)}X_{1}^{n}[\mathcal{G}_{2}\cup\mathcal{B}_{2}];Z_{2}|X_{1}^{n}[\mathcal{G}_{1}\cup\mathcal{B}_{1}]S_{1}^{(1)}S_{2}^{(1)})$
    $=I(S_{1}^{(2)}S_{2}^{(2)}X_{1}^{n}[\mathcal{G}_{2}\cup\mathcal{B}_{2}];Z_{2}).$
@@ IV-A3 Privacy of the unselected strings for the senders
We now prove that the client does not learn any information about the unselected files in (10)
$I\left(Z_{1}Z_{2}Y^{n}\mathbf{A};F_{1,\bar{Z}_{1}}F_{2,\bar{Z}_{2}}\right)$ $\stackrel{{\scriptstyle(a)}}{{=}}I\left(Z_{1}Z_{2}Y^{n}M_{11}M_{12}M_{21}M_{22};F_{1,\bar{Z}_{1}}F_{2,\bar{Z}_{2}}\right)$
$\stackrel{{\scriptstyle(b)}}{{=}}\frac{1}{4}\sum_{z_{1},z_{2}}I\left(Y^{n}M_{11}M_{12}M_{21}M_{22};F_{1,\bar{z}_{1}}F_{2,\bar{z}_{2}}|Z_{1}=z_{1},Z_{2}=z_{2}\right)$
$\stackrel{{\scriptstyle(c)}}{{=}}\frac{1}{4}\sum_{z_{1},z_{2}}I\left(Y^{n}F_{1,z_{1}}F_{2,z_{2}},X_{1}^{n}[S^{(1)}_{\bar{z}_{1}}]\oplus F_{1,\bar{z}_{1}},X_{2}^{n}[S^{(2)}_{\bar{z}_{2}}]\oplus F_{2,\bar{z}_{2}};F_{1,\bar{z}_{1}}F_{2,\bar{z}_{2}}|Z_{1}=z_{1},Z_{2}=z_{2}\right)$
$\stackrel{{\scriptstyle(d)}}{{=}}\frac{1}{4}\sum_{z_{1},z_{2}}I\left(Y^{n},X_{1}^{n}[S^{(1)}_{\bar{z}_{1}}]\oplus F_{1,\bar{z}_{1}},X_{2}^{n}[S^{(2)}_{\bar{z}_{2}}]\oplus F_{2,\bar{z}_{2}};F_{1,\bar{z}_{1}}F_{2,\bar{z}_{2}}|Z_{1}=z_{1},Z_{2}=z_{2}\right)$
$\stackrel{{\scriptstyle(e)}}{{=}}0$ (10)
@@
[…]
@@ IV-A4 Servers’ mutual privacy
$I\left(F_{1,1}F_{1,2}X_{1}^{n}\mathbf{A};F_{2,1}F_{2,2}\right)$
$\stackrel{{\scriptstyle(a)}}{{=}}I(F_{1,1}F_{1,2}X_{1}^{n}M_{21}M_{22}S_{1}^{(1)}S_{2}^{(1)}S_{1}^{(2)}S_{2}^{(2)};F_{2,1}F_{2,2})$
$\stackrel{{\scriptstyle(b)}}{{=}}0,$
@@ IV-B Case $L_{1}=L_{2}$
In this section, we present an achievability scheme for the case $L_{1}=L_{2}=L$ separately, as it is significantly simpler than the achievability scheme for the general case $L_{1}\neq L_{2}$, which is discussed in Section IV-C. The idea is to construct a coding scheme for this case by utilizing L1L-1 times Algorithm 2 developed for the case $L_{1}=L_{2}=2$. This reduction idea is inspired by [brassard1996oblivious] in the context of oblivious transfer. Before we describe our coding scheme in Algorithm 2, we illustrate the coding scheme idea in the following example.
@@
Suppose that $L_{1}=L_{2}=3$, and $S$ and $T$ are uniformly distributed over $\llbracket 1,2^{nR_{1}}\rrbracket$ and $\llbracket 1,2^{nR_{2}}\rrbracket$, respectively. Servers 1 and 2 engage in a dual-source SPIR protocol from Section IV-A, interacting with the client twice. In the first interaction, the client selects between $S$ and $F_{1,1}$ from Server 1, and between $T$ and $F_{2,1}$ from Server 2. During the second interaction, the client chooses between $F_{1,2}\oplus S$ and $F_{1,3}\oplus S$ from Server 1, and between $F_{2,2}\oplus T$ and $F_{2,3}\oplus T$ from Server 2. Specifically, in the first round, if the client desires $F_{1,1}$, they request $F_{1,1}$ from Server 1, otherwise, they request $S$. In the second round, the client requests $F_{1,2}\oplus S$ or $F_{1,3}\oplus S$ from Server 1, depending on whether they wish to retrieve $F_{1,2}$ or $F_{1,3}$, respectively. The same logic applies to Server 2. More specifically, the series of requests made by the client to retrieve the desired files is summarized in Table II.
TABLE II: Series of requests for file selection at each server in Example 1.
Server 1 Server 2
Request 1 $F_{1,1}$ $S$ $S$ $F_{2,1}$ $T$ $T$
Request 2 $F_{1,2}\oplus S$ $F_{1,2}\oplus S$ $F_{1,3}\oplus S$ $F_{2,2}\oplus T$ $F_{2,2}\oplus T$ $F_{2,3}\oplus T$
Recover $F_{1,1}$ $F_{1,2}$ $F_{1,3}$ $F_{2,1}$ $F_{2,2}$ $F_{2,3}$
Algorithm 2 extends the idea of Example 1 to $L_{1}=L_{2}=L$, using L1L-1 sequential interactions between the server and the client. At Server $i\in\{1,2\}$, each file, except the first, is hidden with random masks $(S_{i,t})$, for $t\in\llbracket 1,L-2\rrbracket$. In each round $t\in\llbracket 1,L-1\rrbracket$, the client selects one of two values $(C_{i,t}[1],C_{i,t}[2])$, which are defined as some XOR combinations of masks and files, from Server $i$ via Algorithm 1. Across the L1L-1 rounds, the client carefully chooses which values to recover, accumulates them, and combines them with XOR operations to reconstruct the desired file $F_{i,Z_{i}}$ from Server $i$.
Algorithm 2 Dual-source SPIR when $L_{1}=L_{2}=L$
1:L2L-2 sequences $(S_{1,t})_{t\in\llbracket 1,L-2\rrbracket}$ uniformly distributed over $\{0,1\}^{nR_{1}}$, L2L-2 sequences $(S_{2,t})_{t\in\llbracket 1,L-2\rrbracket}$ uniformly distributed over $\{0,1\}^{nR_{2}}$, the file selection $(Z_{1},Z_{2})\in\mathcal{L}_{1}\times\mathcal{L}_{2}$
@@
$(C_{j,t}[1],C_{j,t}[2])$ $\!\triangleq\!(F_{j,t}\oplus S_{j,t-1},S_{j,t-1}\oplus S_{j,t})$
@@
where $t\in\llbracket 2,L-2\rrbracket$. For $t\in\llbracket 1,L-1\rrbracket$, define $C_{j,t}\triangleq(C_{j,t}[1],C_{j,t}[2])$.
@@
5:  The client and the two servers perform the SPIR protocol in Algorithm 1 with the two sequences $(C_{j,t}[1],C_{j,t}[2])$ at Server $j\in\llbracket 1,2\rrbracket$ and the selection $(Z_{1,t},Z_{2,t})$ for the client.
@@
From Section IV-A5, the lengths of the files in Algorithm 1 for the case $L_{1}=L_{2}=2$ and $n$ channel uses is $\left(\alpha M,\bar{\alpha}M\right)$, hence, the rate pair achieved with Algorithm 2, which utilizes L1L-1 times Algorithm 1 is